OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i

OpenStudy (anonymous):

A: f(x) = x^4 - 362.5x^2 + 1450x - 4984 B: f(x) = x^4 - 9x^3 + 32x^2 - 725x + 4984 C: f(x) = x^4 - 67x^2 + 1450x - 4984 D: f(x) = x^4 - 9x^3 - 32x^2 + 725x - 4984

OpenStudy (anonymous):

Thank you! i'm sprry but I dont understand due to so many of these %%% symbols that interfere with your answer.

OpenStudy (amistre64):

thats what happens when you copy and paste material instead of taking the time to actually work thru the questions with the asker :/

OpenStudy (amistre64):

we should prolly start with setting up the product of the roots

OpenStudy (amistre64):

for any given root "r"; the term (x-r) needs to be included in the setup

OpenStudy (anonymous):

Do i multipy (x-4) (x+14) and something??

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

idk the last thing i multiply it with

OpenStudy (amistre64):

the complex number has a conjugate as well

OpenStudy (anonymous):

is it 5 - 8i

OpenStudy (amistre64):

complex roots always come in pairs 5 + 8i and 5 - 8i so include in the rooted setup (x-(5+8i)) and (x-(5-8i))

OpenStudy (anonymous):

so i multiply those two and (x-14) (x+4)?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (amistre64):

youre welcome, and good luck ;)